OpenStudy (anonymous):

Integral help! I'm just not seeing how to start to solve this integral.....its the integral of (ln(x+3))/(x+3)

OpenStudy (freckles):

did you try letting u=ln(x+3)

OpenStudy (anonymous):

then what would make the (x+3) on the bottom?

OpenStudy (freckles):

\[\int\limits_{}^{}\ln(x+3) \cdot \frac{1}{x+3} dx\] \[u=\ln(x+3) => du=\frac{1}{x+3} dx\]

OpenStudy (freckles):

Now just make those replacements and you will have something easy to integrate

OpenStudy (anonymous):

oh! herp derp to myself

OpenStudy (freckles):

:) I'm glad I could be a service